# Coupled Laplacian Eigenmaps for Locally-Aware 3D Rigid Point Cloud Matching

### v1

@@ 4.1 Experiment settings
L
R
L
R
L
R
L
R
Human
Fisher et al. [19]
18
19
20
20
SSM-Tibia [32]
30
30
ICL [47]
35
35
35
35
35
35
@@
Table 2: Anomaly localization results. The area under the PRO curve is reported for an integration limit of $0.3$ for each evaluated method and dataset category. GAN, AE and VM results are provided by Bergmann et al. [7], and, for a fair comparison, we include the results obtained by restricted GPS [55], eigenfunctions (Hist) [42, 60] and Euclidean matching. The subsctipt on a method indicates the number of used eigenmaps. The overall best performing methods are highlighted in boldface, while the bests for each category are underlined.
Method
Bagel
Cable Gland
Carrot
Cookie
Dowel
Foam
Peach
Potato
Rope
Tire
Mean
3D RGB
VM
0.388
0.321
0.194
0.570
0.408
0.282
0.244
0.349
0.268
0.331
0.335
GAN
0.421
0.422
0.778
0.696
0.494
0.252
0.285 0.362
0.402
0.631
0.474
AE 0.432
0.158
0.808
0.491
0.841 0.406
0.262
0.216
0.716
0.478
0.481
3D Only
GAN
0.111
0.072
0.212
0.174
0.160
0.128
0.003
0.042
0.446
0.075
0.143
AE
0.147
0.069
0.293
0.217
0.207
0.181
0.164
0.066
0.545
0.142
0.203
VM
0.280
0.374
0.243
0.526
0.485
0.314
0.199
0.388
0.543
0.385
0.374
[…]
0.452
0.616
0.695
0.364
0.738
0.471
0.659
0.844
0.647
0.651
0.613
[…]
0.465
0.621
0.690
0.363
0.739
0.480
0.672
0.833
0.653
0.670
0.619
[…]
0.476
0.629
0.703
0.365
0.744
0.473
0.661
0.840
0.647
0.693
0.623
[…]
0.491
0.629
0.698
0.351
0.746
0.501
0.663
0.841
0.652
0.695
0.627
Euclidean
0.655
0.631
0.743
0.615
0.803
0.528
0.726
0.875
0.762
0.695
0.703
$\textbf{Ours}_{100}$
0.669
0.642 0.808
0.714
0.812
0.582
0.748
0.897
0.750
0.733
0.736
$\textbf{Ours}_{200}$ 0.702
0.630
0.728
0.735 0.812 0.701 0.780 0.914 0.767
0.713
0.748
Table 3: Accuracy of human BSE for each bone structures. All the matching methods are applied after RANSAC registration. The overall best performing methods are highlighted in boldface.
Method
Femur
Hip
Fibula
Tibia
Mean
Hausdorff
73.52
97.95
59.30
72.52
75.82
Chamfer
71.64
98.65
64.62
74.47
77.35
FPFH [57]
68.88
96.67
66.69
78.35
77.65
$\textbf{Ours}_{20}$
78.79
98.78
71.28
78.46
81.83
$\textbf{Ours}_{10}$ 79.68
97.76
73.47 78.66 82.39
Bone Side Estimation. The proposed global shape matching score is used to perform automatic BSE, that is, given a reference bone of known side we want to assess the side of another bone, referred as target. Assuming general unknown initial frames, we used the Principal Component Analysis (PCA) to generate a mirrored version of the reference surface. Namely, we identified that the bilateral symmetry of human and animal bodies is equivalent to a mirroring around the second principal component on the vast majority of bones, such as the ones considered in this study. Hence, given the three point clouds, target, source and mirrored source, the Coupled Laplacian of Fig. 2 can be applied. More specifically, we selected the Random Sample Consensus (RANSAC) [20], preceded by a spectral scaling, as rigid registration method. The scaling is performed using only the information carried by the Fiedeler vector, i.e. the eigenvector $\boldsymbol{\phi}_{1}$ corresponding to the first non-zero eigenvalue $\lambda_{1}$. A rough estimation of the bones length, independently of the Euclidean frame, is given by the distance between the points corresponding to minimum and maximum values of the Fiedler vectors. Thanks to that, one of the two bones can be scaled in order to match the length of the other and improve the RANSAC registration. The three aligned $m$-dimensional embeddings, derived from the eigendecomposition of the Coupled Laplacian, are then used as in Eq. 5 to retrieve the best matching source and, consequently, the target side. Note that, the proposed BSE method is fully independent of the target and reference frames and therefore can be applied directly on the segmentation obtained from a medical image, without any previous knowledge, in a fully automated pipeline. Additionally, in alternative, one can using two different bone references, one left and one right, avoiding the mirroring step, or apply the mirroring on the target shape while letting the source unchanged. Full details of the described algorithm are provided in the Supplementary Material.
We generated a benchmark to test the BSE task by collecting several bone surface point cloud data from public datasets [19, 32, 47]. Namely, from Fisher et al. [19] we extracted Femur and Hip structures, from SSM-Tibia [32] only right Tibia and Fibula, and from the Imperial College London (ICL) [47] we collected Femur, Tibia and Fibula point clouds. The composition of the benchmark is shown in Tab. 1 and the detailed instructions to collect and prepare it for testing are provided in the Supplementary Code (or GitHub). Such a variety of shapes, which are acquired using different methods, makes the task more challenging, including some intra-clinic variability. Some additional internal data [40] with knee structures of sheep are used in the Supplementary Material to test the human to animal generalization, and vice-versa, of the proposed BSE technique. In our experiments, for a given bone class, we selected each shape once as source and the others as targets on which the side has to be inferred. Therefore, the average accuracy of all the experiments takes into account the variability of the source shape, and the method robustness can be analysed.
@@
Anomaly Localization. We tested the proposed approach for the 3D anomaly detection task in the MVTec 3D-AD dataset [7]. This recent dataset was designed for the unsupervised detection of anomalies in point clouds of industrially manufactured products. It contains over 4000 high-resolution 3D scans of 10 object categories. In our experiments, we chose one anomaly-free training sample for each class as source and we used the point-wise distances of the aligned eigenmaps between this shape and each target in the test set as anomaly scores. In this case, we performed the registration using affine Coherent Point Drift (CPD) [45], which is more suited for objects like the ones included in the MVTec 3D-AD, e.g. the length and thickness variation among potatoes and carrots samples is better captured by an affine registration rather than rigid. In order to speed-up the computation, we pre-processed the point clouds by removing their flat background with a threshold on the $z$-axis after a 3-dimensional PCA and, without losing generality, we sub-sampled the number of foreground points to a maximum of $13000$. Finally, to compare the obtained anomaly scores with the Ground Truth (GT) image, we projected back the points to the original 2D plane and, if any sub-sampling was performed in pre-processing, we applied a dilation with a structuring element of size the inverse of the sampling factor.
@@ 4.2 Experiment results
We show here comparative quantitative and qualitative results of the proposed method, while a study of the robustness with respect to the source shape and more graphical comparisons are provided in the Supplementary Material.
@@
Bone Side Estimation. In Fig. 3, the average accuracy of BSE is depicted based on the fraction of artificial cross-edges added to compute the Coupled Laplacian and the dimension of the spectral embeddings, $l$ and $m$, respectively. We can observe that the more coupled eigenmaps we aim to utilize for computing a global similarity score, the greater the need for additional cross-connections to ensure the reliability of the results. This occurs because, when graphs are weakly coupled with only a few cross-connections, the Coupled Laplacian captures limited intra-shape characteristics and predominantly emphasizes individual local geometries. In fact, when $l=0$ and therefore the coupled graph is not fully connected, the eigenmaps obtained using the Coupled Laplacian are equivalent to the ones computed independently on each single graph, if the eigenvalues are the same (see Supplmentary Material). On the other hand, not too many cross-edges, nor eigenmaps, are needed to obtain meaningful global matching scores and achieve good performances in the BSE task. In Tab. 3 a quantitative comparison of human BSE with respect to other distance measures is reported. For a fair compariosn, all the other metrics are computed after RANSAC registration of source and mirrored source. Hausdorff and Chamfer discrepancies are calculated on the Euclidean coordinates of the points, while FPFH [57] are 33-dimensional vectors describing each point based on histograms of the surface normals. Our method, using $10$ and $20$ eigenmodes with $l=0.5$ fraction of cross-connections, outperforms the other techniques achieving better accuracy both on the single bones and in average. This is because the description provided by the aligned eigenmaps, obtained through Coupled Laplacian, is more aware of local details than the other distance metrics.
@@
Anomaly Localization. Tab. 2 lists quantitative results for each evaluated method for the localization of anomalies. For each category, the normalized area under the Per-Region Overlap (PRO) curve [7] with an upper integration limit of $0.3$, as well as, the mean performance, are reported. Performance of Generative Adversarial Network (GAN), Autoencoder (AE) and Variation Model (VM) on the same test set are provided by the dataset authors [7]. Furthermore, we tested restricted GPS [55], eigenmaps histogram matching [42, 60] and our method, all with the same pre-processing and source shapes, using $100$ and $200$ eigenmaps. In order to obtain denser and more precise anomaly localization, the Coupled Laplacian is built using the whole set of points, after pre-processing, as cross-connections, i.e. $l=1$. Furthermore, we include the result obtained using as anomaly score for each target point the normalized Euclidean distance of the nearest point in the source geometry after registration. Our method, using only 3D information, outperforms all the other techniques in average and on most of the individual objects. Moreover, we obtained better results than deep-learning methods having RGB and Depth in combination as input. Interestingly, the point-wise similarity computed on eigenmaps not properly aligned (GPS and Hist) is worst than just considering Euclidean distances between points of registered shapes, making worthless the computation of spectral embeddings. In Fig. 4 we compare the qualitative results obtained using different numbers of eigenmaps to score the point-wise similarities. Using a smaller $m$, the proposed technique is prone to individuate only small regions with highly dissimilar local geometries. By increasing the size of the spectral embeddings, we can identify larger and more subtle surface differences, even if they are not necessarily classified as anomalies in the GT. For instance, the potato and bagel surfaces have some natural irregularities, with respect to the source, that are not highlighted using $100$ and $200$-dimensional embeddings, but are instead detected with $300$ eigenmaps. This concept is linked to the modal length, which we define in the Supplementary Material. The trade-off between number of maps used and extent of surface differences detected is interesting to tune the proposed method to other tasks requiring specific attention to identify surface dissimilarities.
@@
Figure 5: Comparison of the distance between source and target, in the 3D anomaly detection task, using the Coupled Laplacian technique and Euclidean distance after RANSAC rigid alignment.
@@
Ablation Study. To further motivate the use of the Laplacian spectral space over the Euclidean space, in Fig. 5 we show some comparison of the distance between source and target, after RANSAC rigid alignment, in the two spaces, the latter obtained using the proposed Coupled Laplacian operator with $m=200$. In the first and second row the better resistance to noise and outliers in the spectral space is highlighted. The bending of the cable gland, i.e. the anomaly, is not detected in the Euclidean space because of an additional segment of points on the back of the target having the largest distance to the source. On the other hand, with the aligned eigenmaps comparison, the anomaly, as well as the noise, is correctly localized. This is because both are local differences, with respect to reference, and, therefore, they can be easily captured in the Laplacian space. In addition, in the third row, we show that our technique provides a sharper anomaly split of the cookie point cloud, which simplifies the anomaly detection task. Nevertheless, the usage of the proposed landmarks-free Coupled Laplacian is limited to tasks in which a correspondence between target and source shapes can be defined by mean of a registration process. Problems including graph isomorphism, such as articulated shape matching [42, 60], or non-rigid shape matching, can be still solved using our method, without relying on the rigid alignment, if some cross-connection are defined a-priori. In this case, landmarks and their neighbouring points would be most likely the easiest option to connect separate graphs and obtain aligned spectral embeddings for further processing. Moreover, a wrong registration of the shapes, due to randomness and errors in the selected registration algorithm, may affect the alignment of the $m$-dimensional embeddings derived from the Coupled Laplacian. Even if the probability of such event for similar geometries is relatively low and the adjacency matrix is already weighted with RBF, a double-check can be added in the pipeline to verify the correct registration [13].
@@ 5 Conclusions
We presented a versatile method to perform 3D rigid point cloud matching, both globally and locally, by using aligned eigenspaces for two or more similar shapes without prior knowledge on the local frames or specif markers on them. The introduction of the Coupled Laplacian operator enables the generation of aligned eigenmaps relying on cross-connections added between graphs of registered geometries. We introduced a novel task consisting in the automatic detection of a bone side, i.e. Bone Side Estimation (BSE), and we proposed a benchmark to test it. Leveraging global similarities of eigenmaps derived from the Coupled Laplacian, we define a full pipeline to perform BSE on arbitrary bone surfaces segmented from a generic medical image. Moreover, we tested the ability of the proposed method in capturing surface differences by performing local shape matching for the 3D anomaly detection task on the MVTec 3D-AD dataset. The proposed technique outperforms other methods on the two tasks, therefore capturing better both global and local similarities, thanks to the matching performed on correctly aligned spectral spaces. Beside the applications showcased here, we believe that our work can help several tasks in many fields in which the localization of local difference is crucial for matching, detection or retrieval. Future extensions could explore generalizations of the technique, not only for rigid matching but also to address challenges such as graph isomorphism. This approach opens new avenues for advancing the State-of-The-Art in 3D shape analysis and matching techniques.

### v2

@@ 4.1 Experiment settings
L R L R L R L R
Human Fisher et al. [18] 18 19 20 20 - - - -
SSM-Tibia [31] - - - - - 30 - 30
ICL [47] 35 35 - - 35 35 35 35
@@
Table 2: Anomaly localization results. The area under the PRO curve is reported for an integration limit of $0.3$ for each evaluated method and dataset category. GAN, AE and VM results are provided by Bergmann et al. [7]. Moreover, we include the results obtained by restricted GPS [55], eigenfunctions (Hist) [42, 60] and Euclidean matching and MAP [39]. All the matching methods are applied after affine CPD registration, when not specified, or CMM + CPD Non-Rigid (NR) registration. The subsctipt on a method indicates the number of eigenmaps used. The overall best performing methods are highlighted in boldface, while the bests for each category are underlined.
Method Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Mean $\uparrow$
3D RGB VM 0.388 0.321 0.194 0.570 0.408 0.282 0.244 0.349 0.268 0.331 0.335
GAN 0.421 0.422 0.778 0.696 0.494 0.252 0.285 0.362 0.402 0.631 0.474
AE 0.432 0.158 0.808 0.491 0.841 0.406 0.262 0.216 0.716 0.478 0.481
3D Only GAN 0.111 0.072 0.212 0.174 0.160 0.128 0.003 0.042 0.446 0.075 0.143
AE 0.147 0.069 0.293 0.217 0.207 0.181 0.164 0.066 0.545 0.142 0.203
VM 0.280 0.374 0.243 0.526 0.485 0.314 0.199 0.388 0.543 0.385 0.374
Euclidean (NR) 0.404 0.623 0.731 0.366 0.771 0.303 0.590 0.772 0.697 0.583 0.584
GPS100[55] 0.452 0.616 0.695 0.364 0.738 0.471 0.659 0.844 0.647 0.651 0.613
GPS200[55] 0.465 0.621 0.690 0.363 0.739 0.480 0.672 0.833 0.653 0.670 0.619
Hist100 [60] 0.476 0.629 0.703 0.365 0.744 0.473 0.661 0.840 0.647 0.693 0.623
MAP200 [39] 0.481 0.630 0.694 0.399 0.742 0.497 0.653 0.832 0.649 0.675 0.625
Hist200 [60] 0.491 0.629 0.698 0.351 0.746 0.501 0.663 0.841 0.652 0.695 0.627
Euclidean 0.655 0.631 0.743 0.615 0.803 0.528 0.726 0.875 0.762 0.695 0.703
$\textbf{Ours}_{100}$ 0.669 0.642 0.808 0.714 0.812 0.582 0.748 0.897 0.750 0.733 0.736
$\textbf{Ours}_{200}$ 0.702 0.630 0.728 0.735 0.812 0.701 0.780 0.914 0.767 0.713 0.748
Table 3: Accuracy [%] of human BSE. All the matching methods are applied after RANSAC registration with spectral scaling, when not specified, or CMM + CPD Non-Rigid (NR) registration. The overall best performing methods are highlighted in boldface.
Method Femur Hip Fibula Tibia Mean $\uparrow$
Chamfer (NR) 56.58 86.58 57.89 74.74 68.95
Hausdorff (NR) 57.37 87.63 59.47 73.68 69.54
Hausdorff 73.52 97.95 59.30 72.52 75.82
Chamfer 71.64 98.65 64.62 74.47 77.35
FPFH [57] 68.88 96.67 66.69 78.35 77.65
MAP20 [39] 76.05 98.54 71.32 74.21 80.13
Hist20 [60] 77.63 98.42 69.47 77.37 80.72
$\textbf{Ours}_{20}$ 78.79 98.78 71.28 78.46 81.83
$\textbf{Ours}_{10}$ 79.68 97.76 73.47 78.66 82.39
Bone Side Estimation. The proposed global shape matching score is used to perform automatic BSE. Assuming general unknown initial frames, we used the Principal Component Analysis (PCA) to generate a mirrored version of the reference surface, i.e. a synthetic contralateral bone. Namely, we identified that the bilateral symmetry of human and animal bodies is equivalent to a mirroring around the second principal component on the vast majority of bones, such as the ones considered in this study. Hence, given the three point clouds, target, source and mirrored source, the Coupled Laplacian of Fig. 2 can be applied. More specifically, we selected the Random Sample Consensus (RANSAC) [19], preceded by a spectral scaling, as rigid registration method to handle varying bone lengths without altering local structures. The scaling is performed using only the information carried by the Fiedeler vector, i.e. the eigenvector $\boldsymbol{\phi}_{1}$ corresponding to the first non-zero eigenvalue $\lambda_{1}$. A rough estimation of the bones length, independently of the Euclidean frame, is given by the distance between the points corresponding to minimum and maximum values of the Fiedler vectors. Thanks to that, one of the two bones can be scaled in order to match the length of the other and improve the RANSAC registration. The three aligned $m$-dimensional embeddings, derived from the eigendecomposition of the Coupled Laplacian, are then used as in Eq. 5 to retrieve the best matching source and, consequently, the target side. Note that, the proposed BSE method is fully independent of the target and reference frames and therefore can be applied directly on the segmentation obtained from a medical image, without any previous knowledge, in a fully automated pipeline. In alternative, one can using two different bone references, one left and one right, avoiding the mirroring step, or apply the mirroring on the target shape while letting the source unchanged. Full details of the described algorithm are provided in the Supplementary Material.
We generated a benchmark to test the BSE task by collecting several bone surface point cloud data from public datasets [18, 31, 47]. Namely, from Fisher et al. [18] we extracted Femur and Hip structures, from SSM-Tibia [31] only right Tibia and Fibula, and from the Imperial College London (ICL) [47] we collected Femur, Tibia and Fibula point clouds. The detailed composition of the benchmark is reported in Tab. 1. Such a variety of shapes, which are acquired using different methods, makes the task more challenging, including some intra-clinic variability. In our experiments, for a given bone class, we performed cross-testing by selecting each shape once as source and the others as targets on which the side has to be inferred. Therefore, the average accuracy of all the experiments, achieved when predicting the correct body side, i.e. binary classification, takes into account the variability of the source shape. The robustness of the method with respect to the reference is discussed in the Supplementary Material.
@@
Anomaly Localization. We tested the proposed approach for the 3D anomaly detection task in the MVTec 3D-AD dataset [7]. This recent dataset was designed for the unsupervised detection of anomalies in point clouds of industrially manufactured products. It contains over 4000 high-resolution 3D scans of 10 object categories. In our experiments, we chose one anomaly-free training sample for each class as source and we used the point-wise distances of the aligned eigenmaps between this shape and each target in the test set as anomaly scores. We performed the pre-registration using affine Coherent Point Drift (CPD) [45], which is more suited for objects like the ones included in the MVTec 3D-AD, e.g. length and thickness variations are better captured by an affine registration rather than rigid. In order to speed-up the computation, we pre-processed the point clouds by removing their flat backgrounds with a threshold on the $z$-axis after a 3-dimensional PCA and, without losing generality, we sub-sampled the number of foreground points to a maximum of $13000$. Finally, to compare the obtained anomaly scores with the Ground Truth (GT) image, we projected back the points to the original 2D plane and, if any sub-sampling was performed, we applied a dilation with a structuring element of size the inverse of the sampling factor.
@@ 4.2 Experiment results
Bone Side Estimation. In Fig. 3, the average accuracy of BSE is depicted based on the fraction of artificial cross-edges added to create the coupled graph and the dimension of the spectral embeddings, $l$ and $m$, respectively. We can observe that the more coupled eigenmaps we aim to utilize for computing a global similarity score, the greater the need for additional cross-connections to ensure the reliability of the results. This occurs because, when graphs are weakly coupled with only a few cross-connections, the Coupled Laplacian captures limited intra-shape characteristics and predominantly emphasizes individual local geometries. In fact, when $l=0$ and therefore the coupled graph is not fully connected, the eigenmaps obtained using the Coupled Laplacian are equivalent to the ones computed independently on each single graph, if the eigenvalues are the same (proof in the Supplmentary Material). On the other hand, not too many cross-edges, nor eigenmaps, are needed to obtain meaningful global matching scores and achieve good performances in the BSE task. In Tab. 3 a quantitative comparison with other methods of human BSE accuracy is reported. For a fair comparison, all methods are performed after RANSAC registration of the sources. Hausdorff and Chamfer discrepancies are calculated on the Euclidean coordinates of the points, for completeness, also in case of 2-Step non-rigid GMM + CPD registration ($2000$x slower than RANSAC). Our method, using $10$ and $20$ eigenmodes with $l=0.5$ fraction of cross-connections, outperforms the other techniques achieving higher accuracy both on the single bones and in average. Hence, the description provided by the aligned eigenmaps, obtained with Coupled Laplacian, is more aware of local details than the other methods.
@@
Anomaly Localization. Tab. 2 lists quantitative results of each evaluated method for the localization of anomalies. For each category, the normalized area under the Per-Region Overlap (PRO) curve with an upper integration limit of $0.3$ [7], as well as, the mean performance, are reported. Performance of Generative Adversarial Network (GAN), Autoencoder (AE) and Variation Model (VM) on the same test set are provided by the dataset authors [7]. Furthermore, we tested restricted GPS [55], eigenmaps histogram matching [42, 60], MAP [39] and our method, all with the same pre-processing and source shapes, using $100$ and $200$ eigenmaps. In order to obtain denser and more precise anomaly localization, the coupled graph is built using the whole set of points as cross-connections, i.e. $l=1$. Furthermore, we include the result obtained using as anomaly score for each target point the normalized Euclidean distance of the nearest point in the source geometry, both using rigid and non-rigid registration. Our method, using only 3D information, outperforms all the other techniques. Moreover, we obtained better results than deep-learning methods having RGB and Depth in combination as input. Interestingly, the point-wise similarity computed on eigenmaps not properly aligned (GPS and Hist) is worst than just considering Euclidean distances between points of registered shapes, making worthless the computation of spectral embeddings. In Fig. 4 we compare the qualitative results obtained using different numbers of eigenmaps to score the point-wise similarities. Using a smaller $m$, the proposed technique is prone to individuate only small regions with highly dissimilar local geometries. By increasing the size of the spectral embeddings, we can identify larger and more subtle surface differences, even if they are not necessarily classified as anomalies in the GT. For instance, the potato and bagel surfaces have some natural irregularities, with respect to the source, that are not highlighted using $100$ and $200$-dimensional embeddings, but are instead detected with $300$ eigenmaps. This concept is linked to the modal length, which we define in the Supplementary Material. The trade-off between number of maps and extent of differences detected is interesting to tune the method to other tasks requiring specific attention to identify surface dissimilarities.
@@
Figure 5: Comparison of the anomaly score obtained using the Coupled Laplacian technique, with $m=200$ and $l=1$, and Euclidean distance, both after affine CPD alignment.
@@
Ablation Study. To further motivate the preference for Coupled Laplacian spectral space over the Euclidean one, in Fig. 5, we display comparisons of point-wise distances (anomaly scores) between source and target in both spaces, after the same registration. In the first and second row the better resistance to noise and outliers in the spectral space is highlighted. The bending of the cable gland, i.e. the anomaly, is barely detected in the Euclidean space because of noisy points on the back. Instead, with the aligned eigenmaps comparison, the anomaly, as well as the noise, is correctly localized. This is because both are local differences, with respect to the reference, and, therefore, they can be easily captured with coupled eigenmaps. In addition, in the third row, we show that our technique provides a sharper anomaly split of the cookie point cloud, which simplifies the anomaly detection task. Nevertheless, the usage of the proposed landmarks-free Coupled Laplacian is limited to tasks in which a correspondence between target and source shapes can be defined by mean of a registration process. Problems including graph isomorphism, such as articulated shape matching [42, 60], or non-rigid shape matching, can be still solved using our method, without relying on the rigid alignment, if some cross-connection are defined a-priori. In this case, landmarks and their neighbouring points would be most likely the easiest option to connect separate graphs and obtain aligned spectral embeddings for further processing. Moreover, a wrong registration of the shapes, due to randomness and errors in the selected registration algorithm, may affect the alignment of the $m$-dimensional embeddings derived from the Coupled Laplacian. We report here some statistics on registration failures in BSE (failure rate [%] / BSE accuracy on failures [%]), using different references, for femur 6.84 $\pm$ 7.82 / 96.67 $\pm$ 11.30 (high symmetry) and hip 0.0 $\pm$ 0.0 (low symmetry). Moreover, to test robustness, we added white Gaussian rotation and translation noise, i.i.d. in the 3 axis post-registration, obtaining (Rot. SD [] / Trans. SD [mm] / Acc. Drop [%]): 2.5 / 5 / 8.16 - 5 / 5 / 15.13 - 7.5 / 10 / 18.16. Given the low registration failure rates and high robustness to noise, we can rely on the proposed coupling algorithm, making the method independent of the acquisition frames. Moreover, the method can well generalize to any chiral shape matching problem and to tasks like identification of topological noise in graphs and partial matching (see example in Supplementary Material).
@@ 5 Conclusions
We presented a versatile method to perform 3D rigid point cloud matching, both globally and locally, by using aligned eigenspaces for two or more similar shapes without prior knowledge on the local frames or specif markers on them. The introduction of the Coupled Laplacian operator enables the generation of aligned eigenmaps relying on cross-connections added between graphs of registered geometries. We introduced a novel task consisting in the automatic detection of a bone side, i.e. Bone Side Estimation (BSE), and we proposed a benchmark to test it. Leveraging global similarities of eigenmaps derived from the Coupled Laplacian, we define a full pipeline to perform BSE on arbitrary bone surfaces segmented from a generic medical image. Moreover, we tested the ability of the proposed method in capturing local surface differences by performing 3D anomaly detection on the MVTec 3D-AD dataset. The proposed technique outperforms other methods on the two tasks, therefore capturing better both global and local similarities, thanks to the matching performed on correctly aligned spectral spaces. Beside the applications showcased here, we believe that our work can help several tasks in many fields in which the localization of local difference is crucial for matching, detection or retrieval. Future extensions could explore generalizations of the technique, not only for rigid matching but also to address challenges such as graph isomorphism. This approach opens new avenues for advancing the State-of-The-Art in 3D shape analysis and matching techniques.